OpenStudy (anonymous):

8y-4y+10=-y stp by step plz

OpenStudy (anonymous):

8y-4y+10=-y, Notice we can do 8y-4y=4y, on the left hand side Combine like terms first: 4y+10=-y Now subtract 4y from both sides:4y+10-4y=-y-4y Simplify: 10=-5y Divide both side by -5 and get:-2=y

OpenStudy (anonymous):

to cofussing for me

OpenStudy (anonymous):

Okay, 8y-4y+10=-y 4y+10=-y 10=-5y -2=y